OpenStudy (anonymous):

Medal Which exponent makes the statement true? 1/5^9= 5^?

jhonyy9 (jhonyy9):

1/a^2 = a^(-2) this will help you ?

jhonyy9 (jhonyy9):

so this mean 1/5^9 = 5^(-9) hope is understandably